OpenStudy (anonymous):

Imagine you and your partner are fountain designers and you need to know how different speed motors will affect the water’s projected height and time in the air. You have various motor speed options but need to choose the one that most closely meets the client’s height request of 690 feet. Your job is to determine which motor speed is the best choice. You will determine the speed of two of the motors and your partner will do the same for the other two motors. In all cases, you will use the projectile formula y = –16x2 + vx + s, with startin

OpenStudy (anonymous):

I got this Choreographed Fountains Activity: Collaborative Work Imagine you and your partner are fountain designers and you need to know how different speed motors will affect the water’s projected height and time in the air. You have various motor speed options but need to choose the one that most closely meets the client’s height request of 690 feet. Your job is to determine which motor speed is the best choice. You will determine the speed of two of the motors and your partner will do the same for the other two motors. In all cases, you will use the projectile formula y = –16x2 + vx + s, with starting height of 0 (s = 0). Make sure that you and your partner complete equal amounts of this activity. 1. Substitute the data provided for speed (v) and starting heights (s) into their proper positions in the formula y = –16x2 + vx + s and write the equation corresponding with each motor in the space provided below. Then identify the a, b, and c values for each where indicated. (4 points total, 1 point each) The equation for motor A is: y = –16x2 + 190x + 0 a =-16 b =190 c =690 The equation for motor B is: y = –16x2 + 210x + 0 a = -16 b = 210 c = 690 The equation for motor C is: y = –16x2 + 210x + 0 a =-16 b =230 c =690 The equation for motor D is: y = –16x2 + 210x + 0 a =-16 b = 250 c =690

OpenStudy (anonymous):

I don't get how to graph it I entered the formula into the input and it's just two lines

OpenStudy (anonymous):

610 = –16x^2 + 190x plot 610= –16x^2 + 210x plot 610= –16x^2 + 250x plot

OpenStudy (anonymous):

are those suppose to be just lines

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

another girl i helped on yahoo had some of the same problems look at this

OpenStudy (anonymous):

oh how do I find the vertex though

OpenStudy (anonymous):

heres how you find the vertex

OpenStudy (anonymous):

you type in y = -16x^2 + 230x for motor C

OpenStudy (anonymous):

http://www.purplemath.com/modules/sqrvertx.htm

OpenStudy (anonymous):

I got this graph

OpenStudy (anonymous):

and y = -16x^2 + 250x for motor D

OpenStudy (anonymous):

woah... um let me c im not sure i have to do some research